A sum of money is shared into 3 parts 2:4:5. The largest share is $40. What is the total amount of money shared
Answer by stanbon(75887)

A sum of money is shared into 3 parts 2:4:5. The largest share is $40. What is the total amount of money shared
-----
2:4:5 is the same as 2x:4x:5x
------
Equation:
5x = 40
x = 8
----
Total amount = 2x + 4x + 5x = 11x = 11*8 = 88
==================================================
Cheers,
Stan H.
=============